OpenStudy (anonymous):

Given the parent functions f(x) = 5x − 1 and g(x) = 3x − 9, what is g(x) − f(x)? Please help will give medal

OpenStudy (anonymous):

g(x) − f(x) = 3x − 8 − 5x g(x) − f(x) = 3x − 10 − 5x g(x) − f(x) = −2x − 8 − 3x g(x) − f(x) = −2x − 10 − 3x

OpenStudy (anonymous):

g(x)-f(x)= (3x-9)-(5x-1)= 3x-8-5x

OpenStudy (anonymous):

Your a lifesaver thank you so much, did you just use the foil method? @ninjasandtigers

OpenStudy (anonymous):

Well you didn't have to multiply so no foiling was involved and for some reason the answers aren't completely simplified. But I just combined common factors.

OpenStudy (anonymous):

I see, thank you again!

OpenStudy (anonymous):

you are welcome :)